# A combined Quantum Monte Carlo and DFT study of the strain response and magnetic properties of two-dimensional (2D) 1T-VSe2 with charge density wave

### v1

@@ Abstract
Two-dimensional (2D) 1T-VSe2 has prompted significant interest due to the discrepancies regarding alleged ferromagnetism (FM) at room temperature, charge density wave (CDW) states and the interplay between the two. We employed a combined Diffusion Monte Carlo (DMC) and density functional theory (DFT) approach to accurately investigate the magnetic properties and response of strain of monolayer 1T-VSe2. Our calculations show the delicate competition between various phases, revealing critical insights into the relationship between their energetic and structural properties. We went on to perform Classical Monte Carlo simulations informed by our DMC and DFT results, and found the magnetic transition temperature () of the undistorted (non-CDW) FM phase to be 228 K and the distorted (CDW) phase to be 68 K. Additionally, we studied the response of biaxial strain on the energetic stability and magnetic properties of various phases of 2D 1T-VSe2 and found that small amounts of strain can enhance the , suggesting a promising route for engineering and enhancing magnetic behavior. Finally, we synthesized 1T-VSe2 and performed Raman spectroscopy measurements, which were in close agreement with our calculated results. Our work emphasizes the role of highly accurate DMC methods in advancing the understanding of monolayer 1T-VSe2 and provides a robust framework for future studies of 2D magnetic materials.
@@ I Introduction
Strongly correlated two-dimensional (2D) magnets are a puzzling class of materials from a fundamental physics perspective. With the discovery of ferromagnetism in 2D systems such as CrI3 [1] and Cr2Ge2Te6 [2], an effort to identify and understand the underlying mechanisms of 2D magnets with a finite transition temperature () has become a highly active area of materials science. One of the most interesting and controversial 2D ferromagnetic materials isVSe2, which has a metallic 1T phase (octahedral (1T)-centered honeycombs) and a semiconducting 2H phase (trigonal prismatic (2H)-hexagonal honeycombs) [3]. Despite discrepancies of the structural properties coupled to the energetic stability (whether or not 1T vs. 2H is more favorable) [4, 5, 6, 7] that we were successfully able to resolve in our previous work using highly-accurate electronic structure methods (see Ref. 8), there are several remaining questions regarding 2D 1T-VSe2.
In a 2018 study, single layer 1T-VSe2 was synthesized on van der Waals (vdW) substrates (graphite and MoS2), where strong ferromagnetic ordering was measured above room temperature and a charge density wave (CDW) was detected with a transition temperature of 121 K [9]. This was the first study to demonstrate that the CDW of single layer 1T-VSe2 is coupled with its magnetic properties [9]. Despite room temperature ferromagnetism being reproduced in some cases, such as for a chemically exfoliated monolayer [10], there have been several conflicting reports of the magnetic properties and the interplay between the CDW state and magnetism [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 11, 19, 20, 21, 22]. For example, some studies have found the nonmagnetic CDW phase to be experimentally favorable, with an inherent absence of ferromagnetism [13, 14, 15, 16].
@@
Although our previous semi-local DFT calculations provide a good qualitative assessment of how different magnetic orderings can compete in the undistorted and CDW structures, a more quantitative answer is required. This is required in order to accurately understand the magnetic and CDW transitions in monolayer 1T-VSe2 and estimate quantities such as transition temperatures. For this reason, high-fidelity many-body techniques such as Diffusion Monte Carlo (DMC) [32] can be utilized to accurately describe the electronic and magnetic properties of 2D 1T-VSe2. DMC is a correlated electronic structure method that has a reduced sensitivity on approximations such as the exchange-correlation functional and the Hubbard [33] correction. In addition, DMC has successfully been applied to several 2D and quasi-2D systems [8, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. In this work, we investigate the magnetic properties of monolayer 1T-VSe2 in various magnetic states with and without CDW through the lens of DMC and provide a thorough benchmark using several DFT functionals. In addition to providing accurate quantitative estimates of energy differences between different phases, we coupled our DMC results with classical Monte Carlo simulations to estimate magnetic transition temperatures. Finally, we used our DMC insights to perform DFT calculations of the Raman modes and compared to our own experimental results on a synthesized 1T-VSe2 flake.
II Methodology
Benchmarking DFT calculations for monolayer 1T-VSe2 were performed with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials [50, 51]. For testing purposes, we employed a variety of exchange-correlation functionals including the local density approximation (LDA)[52, 53], Perdew-Burke-Ernzerhof (PBE)[54], the strongly constrained and appropriately normed (SCAN)[55] meta-GGA functional and the r2SCAN [56] meta-GGA functional with, and without the Hubbard correction [33], which was used to treat the on-site Coulomb interaction of orbitals of V. A k-grid of 24x24x1 was used and scaled proportionally to the supercell size, a plane-wave kinetic energy of 400 eV (increased to 800 eV for SCAN and r2SCAN) was used and at least 20 Å of vacuum was added between periodic layers. Our density functional perturbation theory (DFPT) phonon calculations to obtain the peak positions of Raman modes were also performed with VASP along with the phonopy [57, 58] package.
Our Quantum Monte Carlo (QMC) calculations for 1T-VSe2 used the exact same settings as our previous work on 1T- and 2H-VSe2 (see Ref. 8 for more specific details regarding convergence testing). DFT-PBE was used to generate the trial wavefunction (nodal surface) for DMC using the Quantum Espresso (QE) [59] code with a correction of 2 eV (PBE+ = 2 eV resulted in a trial wavefunction with the lowest total energy [8]). We used a k-point grid of 12x12x1 and a kinetic energy cutoff of 4,080 eV (300 Ry) to generate the trial wavefunction. In terms of pseudopotentials, RRKJ (OPT) potentials were used for V [60] and Burkatzki-Fillipi-Dolg (BFD) were used for Se [61]. Variational Monte Carlo (VMC) and DMC [32] simulations were conducted using the QMCPACK [62, 63] code and the Nexus [64] workflow manager was used to automate the DFT-VMC-DMC calculations. Up to two-body Jastrow parameters [65, 66, 67] were optimized using the linear method [68, 69] to minimize the variance and energy in VMC, with the goal of modeling electron correlation and reducing uncertainty in the DMC results [69, 70]. In order to compute the nonlocal part of the pseudopotentials, the locality approximation [70] was used in DMC. An optimal timestep of 0.01 Ha-1 (0.27 eV-1) [8] was used in DMC and to reduce finite-size errors, twist averaging and extrapolating to larger supercells (up to 72 atoms) was performed. A more detailed explanation of the theory behind VMC and DMC can be found in Ref. 32.
Energies of different spin configurations derived from first-principles calculations were mapped onto a classical spin Hamiltonian, characterizing three-dimensional spins arranged on a two-dimensional lattice at sites of V atom. Subsequently, this Hamiltonian served as the basis for executing classical Monte Carlo simulations aimed at estimating the transition transition temperature () for both the undistorted and CDW phases. The Hamiltonian, which has previously been used for other 2D magnetic systems [71, 72, 37], includes the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (),
The indices and iterate over all magnetic sites and their corresponding first nearest neighbor (NN) magnetic sites, respectively. Only the first NN interactions are considered due to the strong localized magnetic moment on V. We have calculated for the CDW phase to be , supporting our assumptions. According to the adopted sign convention, favors isotropic spin alignment, leading to a preferred ferromagnetic (FM) over an antiferromagnetic (AFM) phase. Similarly, favors alignment of the spin -components, and favors the out-of-plane direction as the easy axis. We employ a supercell, equivalent to a nearly nm2 square cell, for the MC simulations. The spin configuration space is discretized into uniformly distributed points on the surface of a unit sphere with a fine resolution of degrees, facilitating detailed exploration of spin orientations. The MC simulation consists of numerous sweeps across all magnetic sites. During each sweep, the spin orientation at each site is updated randomly, one site at a time, and the new configuration is then either accepted or rejected based on the Metropolis algorithm. The convergence of both energy and magnetization is meticulously monitored across varying temperatures. Upon achieving equilibration, an ensemble of configurations is collected to calculate the magnetization and specific heat. We use an ensemble size of k for the CDW simulation and k for the undistorted simulation. Additional methodology details can be found in the SI.
III Results and Discussion
@@
In this work, we specifically focused on monolayer 1T-VSe2 in its freestanding form. Motivated by the experimental discrepancies of controversial room temperature ferromagnetism and competing magnetism and CDW states in the 1T phase, we decided to perform in-depth benchmarking at the DFT level using a variety of approximations. We focused on the normal (undistorted) structures of 2D 1T-VSe2 in its FM, AFM (spins are anti-aligned in a stripy pattern) and nonmagnetic (NM) orientations. In addition to the normal unperturbed crystal structure of 1T-VSe2, we studied the distorted xx1 supercell which is a signature of the material being in its CDW state. This distorted xx1 CDW structure has previously been studied extensively with DFT and experimentally verified. Building on the work of Ref. 26, we performed a thorough DFT benchmark (using several different approaches) of the various magnetic states of the distorted CDW structure. These configurations included FM, NM, and various AFM configurations. Consistent with Ref. 26, we studied the xx1 distorted AFM-A, AFM-B1 and AFM-B2 configurations (see Fig. 1). It is important to note that the distorted AFM-A configuration is equivalent to the AFM (stripy) configuration for the undistorted structure. A summary of these different structures and magnetic configurations are given in Fig. 1.
@@
In our previous work (Ref. 8), we determined the optimal geometry (including lattice parameter and bond distance) and relative phase stability of 2D 1T- and 2H-VSe2 in the FM orientation with DMC and benchmarked with several DFT functionals with and without the Hubbard correction. Fig. 2a) depicts the DFT benchmarking calculations for the NM, FM, and AFM orientations of the undistorted structure. We performed DFT calculations with LDA, PBE, SCAN, r2SCAN with values ranging from (0 to 3) eV, where we fully relaxed each structure in its respective magnetic orientation. As seen in Fig. 2a), which plots the energy differences with respect to the NM state for each functional, the discrepancies with each DFT method are enormous. Although qualitatively the results are somewhat similar (i.e., the FM state being lowest in energy across the board for most DFT functionals), quantitatively, the results vary drastically. In fact, the energy per formula unit (f.u.) between the NM orientation and the FM and AFM orientations can vary up to 0.4 eV. A quantitative energy difference between magnetic states is absolutely necessary to obtain magnetic exchange parameters () and therefore transition temperature () with reasonable accuracy.
@@
In order to overcome these shortcomings of local and semilocal DFT, we performed DMC calculations to obtain accurate total energies of the NM, FM and AFM orientations of the normal undistorted structure of monolayer 1T-VSe2. As a starting geometry for our DMC calculations, we used the structure obtained from our previous work in Ref. 8. We acknowledge that the local magnetic ordering can impact the structural properties, but for our DMC calculations we used the same structure for NM, FM and AFM calculations. In order to study the influence of using the same geometry for NM, AFM and FM on the energetics, we performed DFT calculations with the fixed FM geometry for all magnetic orientations and found a maximum energy difference of 10 meV between the fixed and relaxed structures, which is within the error bar of our DMC calculations. To reduce finite-size errors, we performed DMC calculations for two reasonably sized supercells for NM, FM and AFM and extrapolated to the thermodynamic limit. We used the following supercell sizes: 36 and 72 atoms for NM and AFM, 27 and 48 atoms for FM.The DMC energy differences and uncertainties are depicted in Fig. 2 and Table 1 (in shorthand notation). From the figure and table, we see that SCAN and r2SCAN (with no Hubbard correction) are in closest agreement with our DMC benchmark. For the FM configuration, we also find that SCAN and r2SCAN successfully reproduce the DMC geometry that was previously obtained in Ref. 8.
@@
Although other semilocal DFT functionals such as PBE+ ( = 1, 2 eV) can do a sufficient job at reproducing our DMC geometry, they fail in terms of correctly capturing the quantitative energy differences of various magnetic states. Meta-GGA functionals (SCAN, r2SCAN) can be quite successful for 2D magnetic materials such as VSe2. SCAN and r2SCAN include the kinetic energy density (in addition to the electron density and its gradient) in the exchange-correlation functional, which allows them to capture complicated electronic interactions. These functionals can reduce the self-interaction errors of typical GGAs and improve the description of correlation effects. SCAN satisfies all 17 exact known constraints that a meta-GGA can satisfy, while r2SCAN breaks some of these exact constraints in favor of numerical stability (without significantly hindering accuracy). Although these functionals can simultaneously capture multiple properties such as energy differences between magnetic phases and optimal geometry, SCAN and r2SCAN have been widely reported to overestimate on-site magnetic moments [73, 74, 75, 76]. In the case of 1T-VSe2, we found the DMC on-site magnetic moment of the V atom to be 1.06(2) in Ref. 8. We found the on-site moment to be 1.40 with SCAN and 1.43 with r2SCAN. This discrepancy can be due in part to the tendency of SCAN/r2SCAN to overestimate magnetic moments [73, 74, 75, 76]. It can also be due in part to the limitations of using a single determinant wavefunction in our DMC calculations [77, 78, 79, 80]. Nonetheless, due to the success of multi-property predictions with SCAN and r2SCAN with respect to our DMC benchmark calculations, we decided to use SCAN to obtain more complex quantities (that are difficult to obtain with DMC) such as magnetic anisotropy energies. For computationally demanding simulations such as for the Raman active modes where it is not computationally tractable to use SCAN or r2SCAN, we used PBE+ ( = 1 eV) due to its relative qualitative agreement with DMC results (in terms of structural properties).
@@
Fig. 2c) depicts the results for the CDW distorted structures with various magnetic orientations. The initial supercells for these distorted structures were obtained from Ref. 26, where the V and Se atoms were initially displaced by 0.12 and 0.18 Å respectively (similarly to Ref. 15) prior to atomic relaxation. For our DFT benchmarking calculations, we adapted a slightly different approach for the distorted CDW structures (in contrast to the undistorted cells). For these supercells, we modified the lattice constant to be a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8) and fixed the cell dimensions. We then allowed the atomic positions (with initially displaced atoms) to relax with each respective DFT functional. Similarly to the undistorted structures in Fig. 2a), there is a massive discrepancy in relative energy between each magnetic orientation (with respect to the NM undistorted cell) of the distorted CDW structures (up to 0.5 eV). We observe that the FM orientation is the lowest energy magnetic state for all functionals depicted in Fig. 2b). In contrast to the undistorted structure in Fig. 2a), the energy differences between the FM and various AFM orientations is much smaller. Our results are in good qualitative agreement with the DFT calculations performed in Ref. 26, but are shifted slightly in energy which can be attributed to fixing the lattice constant to the DMC obtained value. Similar to the undistorted case, we also find excellent qualitative agreement between SCAN/r2SCAN and PBE+ ( = 1 eV) for the energy differences between FM and AFM configurations of distorted structures. To provide more insight on these DFT calculations, we performed DMC simulations for just the FM CDW distorted structure at supercells of 30 and 60 atoms and extrapolated to the thermodynamic limit. For the starting geometry of this structure, we optimized the initially displaced atomic coordinates with r2SCAN and used a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8). This DMC energy difference and uncertainty is shown (with respect to the undistorted NM structure) in 2c) and Table 1. We find that the DMC energy difference lies between PBE+ ( = 1 eV) and PBE+ ( = 2 eV). Additionally, we find that the SCAN/r2SCAN and LDA+ ( = 3 eV) energy differences are close to the DMC value.
In addition to benchmarking the energy differences with DFT and DMC for the distorted CDW structures with various magnetic orientations, we decided to study the impact that each DFT functional has on the structural distortions (initially displaced atoms). Interestingly, we found that when the atomic coordinates of some of the distorted CDW structures (with specific DFT functionals and Hubbard corrections) are relaxed, the distortions are significantly reduced and almost disappear. This is most apparent for the distorted AFM-A orientation. For example, when SCAN, r2SCAN, PBE+ ( = 1, 2, 3 eV) and LDA+ ( = 2, 3 eV) are used to relax the AFM-A structure, the distortions disappear completely and the structure relaxes to the normal (undistorted) AFM (stripy) structure. This implies that the stripy AFM ordering cannot exist in the distorted structure. Interestingly, for all DFT functionals considered in this work, the distortions did not disappear from the NM-CDW structure In contrast, we see the distortion disappear in CDW-AFM-B1 for LDA+ and PBE+ ( = 3 eV) and in CDW-AFM-B2 for LDA+ ( = 3 eV). Regarding the CDW-FM structure, we see a significant reduction of the distortion for LDA+ ( = 0, 3 eV) and PBE+ ( = 0, 2, 3 eV). The disappearance of the CDW distortion may be due in part to performing these DFT calculations at a fixed lattice constant predicted by DMC (from Ref. 8) that is different than the lattice constant computed by each respective functional, which can induce strain. The fact that the geometry predicted by SCAN/r2SCAN and PBE+ are closest to the DMC value could be an explanation for why the distortion does not disappear for those functionals (since minimal strain is induced). From these results, we can infer that FM in the distorted-CDW structure is highly sensitive to strain (which is discussed later on in this manuscript). This highlights the competitive nature and interplay of structural parameters, strain mechanisms and magnetic ordering that can occur within the 1T-VSe2 monolayer.
For comparison sake, we analyzed the distortions in the FM-CDW structure obtained from SCAN/r2SCAN (fixed to the scaled DMC lattice constant) to the distortions obtained from PBE calculations in Ref. [26] (where the lattice constant was fixed to 3.36 Å). These comparisons are illustrated in Table S1, where we display the deviation in nearest neighbor bond distances (V-V) with respect to the undistorted bond distances (calculated with SCAN, r2SCAN and PBE respectively) for all of the V-V bonds in the honeycomb lattice (6 bond distances for 5 different V atoms in the CDW supercell). We find that the distortions are reduced in the SCAN/r2SCAN calculations when compared to the previous PBE results [26], which can again be due in part to fixing the lattice constant and inducing strain. Nonetheless, our DMC energies indicate that undistorted FM and CDW-FM are quite close in energy (-0.20(1) eV vs. -0.23(1) eV, with respect to the undistorted NM phase). We propose that the -0.20 eV is coming from the magnetic instability while the -0.03 eV is coming from the CDW, which is consistent with the fact that the magnetic moments (SCAN) between undistorted FM and CDW-FM are nearly identical.
@@
From our DMC insight, we went on to estimate the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (). Since the incorporation of spin-orbit within DMC is a relatively new development and the spin-orbit contributions for 1T-VSe2 are within the DMC uncertainty, we performed collinear DMC calculations (where spins can orient either up or down) for the undistorted FM and AFM orientations (depicted in Fig. 2 and Table 1). For the undistorted structure, we obtain a value of 71(10) meV (no spin-orbit contribution). Taking the insight from our DMC and DFT calculations, where we observed that SCAN reproduced our DMC benchmarking results, we went on to perform spin-orbit SCAN calculations to explicitly obtain anisotropy parameters and (see SI for more detailed information). From these calculations, we obtained a value of -0.77 meV and a value of 0.27 meV. With regards to the distorted-CDW structure, we performed a DMC benchmark calculation for the FM orientation (see Fig. 2 and Table 1). Due to the reasonable agreement with DMC and SCAN and for consistency with the undistorted case, we went on to calculate exchange and anisotropy parameters for the distorted structure using spin-orbit SCAN calculations. In the case of SCAN (since the AFM-A CDW-distorted structure relaxed to the undistorted AFM orientation), the AFM-B2 configuration is the ground state AFM configuration. Using our CDW-FM and CDW-AFM-B2 noncollinear SCAN calculations, we obtained a obtained a value of 21 meV, a value of -0.57 meV, and a value of -0.77 meV (see SI for more details). We went on to use these accurately computed magnetic exchange and anisotropy parameters for our spin Hamiltonian model shown in Eq. 1, where we performed classical Monte Carlo simulations to gain insight on the magnetic phase transitions of the undistorted and distorted structures.

### v2

# Quantum Monte Carlo and density functional theory study of strain and magnetism in 2D 1T-VSe2 with charge density wave states

@@ Abstract
Abstract
Two-dimensional (2D) 1T-VSe2 has prompted significant interest due to the discrepancies regarding alleged ferromagnetism (FM) at room temperature, charge density wave (CDW) states and the interplay between the two. We employed a combined Diffusion Monte Carlo (DMC) and density functional theory (DFT) approach to accurately investigate the magnetic properties, CDW states, and their response to strain in monolayer 1T-VSe2. Our calculations show the delicate competition between various phases, revealing critical insights into the relationship between their energetic and structural properties. We performed classical Monte Carlo simulations informed by our DMC and DFT results, and found the magnetic transition temperature () of the undistorted (non-CDW) FM phase to be 228 K and the distorted (CDW) phase to be 68 K. Additionally, we studied the response of biaxial strain on the energetic stability and magnetic properties of various phases of 2D 1T-VSe2 and found that small amounts of strain can increase the , suggesting a promising route for engineering and enhancing magnetic behavior. Finally, we synthesized 1T-VSe2 and performed Raman spectroscopy measurements, which were in close agreement with our calculated results, validating our computational approach. Our work emphasizes the role of highly accurate DMC methods in advancing the understanding of monolayer 1T-VSe2 and provides a robust framework for future studies of 2D magnetic materials.
Keywords: Quantum Monte Carlo; density functional theory; 2D materials; 2D magnets; strongly correlated materials; strain
@@ I Introduction
Strongly correlated two-dimensional (2D) magnets are a puzzling class of materials from a fundamental physics perspective. With the discovery of ferromagnetism in 2D systems such as CrI3 [1] and Cr2Ge2Te6 [2], an effort to identify and understand the underlying mechanisms of 2D magnets with a finite transition temperature () has become a highly active area of materials science. One of the most interesting and controversial 2D ferromagnetic materials is VSe2, which has a metallic 1T phase (octahedral (1T)-centered honeycombs) and a semiconducting 2H phase (trigonal prismatic (2H)-hexagonal honeycomb)[3]. Despite discrepancies of the structural properties coupled to the energetic stability (whether or not 1T vs. 2H is more favorable) [4, 5, 6, 7] that we were successfully able to resolve in our previous work using highly-accurate electronic structure methods (see Ref. 8), there are several remaining questions regarding 2D 1T-VSe2.
In a 2018 study, single layer 1T-VSe2 was synthesized on van der Waals (vdW) substrates (graphite and MoS2), where strong ferromagnetic ordering was measured above room temperature and a charge density wave (CDW) was detected with a transition temperature of 121 K [9]. This was the first study to demonstrate that the CDW transition of single layer 1T-VSe2 is coupled with its magnetic properties [9]. Despite room temperature ferromagnetism being reproduced in some cases, such as for a chemically exfoliated monolayer [10], there have been several conflicting reports of the magnetic properties and the interplay between the CDW state and magnetism [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 11, 19, 20, 21, 22]. For example, some studies have found the nonmagnetic CDW phase to be experimentally favorable, with an inherent absence of ferromagnetism [13, 14, 15, 16].
@@
Although our previous semi-local DFT calculations [26] provide a good qualitative assessment of how different magnetic orderings can compete in the undistorted and CDW structures, a more quantitative answer is required to accurately understand the magnetic and CDW transitions in monolayer 1T-VSe2 and estimate quantities such as transition temperatures. For this reason, high-fidelity many-body techniques such as fixed-node Diffusion Monte Carlo (DMC) [32] can be utilized to accurately describe the electronic and magnetic properties of 2D 1T-VSe2. DMC is a correlated electronic structure method that has a reduced sensitivity to approximations such as the exchange-correlation functional and the Hubbard [33] correction. In addition, DMC has successfully been applied to several 2D and quasi-2D systems [8, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. In this work, we investigate the magnetic properties of monolayer 1T-VSe2 in various magnetic states and geometries, with and without CDW distortions, through the lens of DMC and provide a thorough benchmark using several DFT functionals. In addition to providing accurate quantitative estimates of energy differences between different phases, we coupled our DMC results with classical Monte Carlo simulations to estimate magnetic transition temperatures. Finally, we used our DMC insights to perform DFT calculations of the Raman modes and compared to our own experimental results on a synthesized 1T-VSe2 flake.
II Results and Discussion
@@
In this work, we specifically focused on monolayer 1T-VSe2 in its freestanding form. Motivated by the experimental discrepancies of controversial room temperature ferromagnetism, competing magnetism and CDW states in the 1T phase, we decided to perform in-depth benchmarking at the DFT level using a variety of approximations. A summary of the different structures and magnetic configurations we studied are given in Fig. 1. We focused on the normal (undistorted) structures of 2D 1T-VSe2 in its FM, AFM (spins are anti-aligned in a stripy pattern) and nonmagnetic (NM) orientations. In addition to the normal unperturbed crystal structure of 1T-VSe2, we studied the distorted xx1 supercell which is a signature of the material being in its CDW state. This distorted xx1 CDW structure has previously been studied extensively with DFT and experimentally verified. Building on the work of Ref. 26, we performed a thorough DFT benchmark (using several different approaches) of the various magnetic states of the distorted CDW structure. These configurations included FM, NM, and various AFM configurations. Consistent with Ref. 26, we studied the xx1 distorted AFM-A, AFM-B1 and AFM-B2 configurations (see Fig. 1). It is important to note that the distorted AFM-A configuration is equivalent to the AFM (stripy) configuration for the undistorted structure.
@@
In our previous work (Ref. 8), we determined the optimal geometry (including lattice parameter and bond distance) and relative phase stability of 2D 1T- and 2H-VSe2 in the FM orientation with DMC and benchmarked with several DFT functionals with and without the Hubbard correction. Fig. S1a) depicts the DFT benchmarking calculations for the NM, FM, and AFM orientations of the undistorted structure. We performed DFT calculations with LDA, PBE, SCAN, r2SCAN with values ranging from (0 to 3) eV, where we fully relaxed each structure in its respective magnetic orientation. As seen in Fig. S1a), which plots the energy differences with respect to the NM state for each functional, the discrepancies with each DFT method are enormous . Although qualitatively the results are somewhat similar (i.e., the FM state being lowest in energy across the board for most DFT functionals), quantitatively, the results vary drastically. In fact, the energy per formula unit (f.u.) between the NM orientation and the FM/AFM orientations can vary up to 0.4 eV. A quantitative energy difference between magnetic states is absolutely necessary to obtain magnetic exchange parameters () and therefore transition temperature () with reasonable accuracy.
@@
In order to overcome these shortcomings of local and semilocal DFT, we performed DMC calculations to obtain accurate total energies of the NM, FM and AFM orientations of the normal undistorted structure of monolayer 1T-VSe2. As a starting geometry for our DMC calculations, we used the structure obtained from our previous work in Ref. 8. We acknowledge that the local magnetic ordering can impact the structural properties, but for our DMC calculations we used the same structure for NM, FM and AFM calculations. In order to study the influence of using the same geometry for NM, AFM and FM on the energetics, we performed DFT calculations with the fixed FM geometry for all magnetic orientations and found a maximum energy difference of 10 meV between the fixed and relaxed structures, which is within the error bar of our DMC calculations. To reduce finite-size errors, we performed DMC calculations for two reasonably sized supercells for NM, FM and AFM and extrapolated to the thermodynamic limit. We used the following supercell sizes: 36 and 72 atoms for NM and AFM, 27 and 48 atoms for FM. The DMC energy differences and uncertainties are depicted in Fig. S1 and Table 1 (in shorthand notation). From the figure and table, we see that SCAN and r2SCAN (with no Hubbard correction) are in closest agreement with our DMC benchmark. For the FM configuration, we also find that SCAN and r2SCAN successfully reproduce the DMC geometry that was previously obtained in Ref. 8.
@@
Although other semilocal DFT functionals such as PBE+ ( = 1, 2 eV) can do a sufficient job at reproducing our DMC geometry, they fail in terms of correctly capturing the quantitative energy differences of various magnetic states. Meta-GGA functionals (SCAN, r2SCAN) can be quite successful for 2D magnetic materials such as VSe2. SCAN and r2SCAN include the kinetic energy density (in addition to the electron density and its gradient) in the exchange-correlation functional, which allows them to capture complicated electronic interactions. These functionals can reduce the self-interaction errors of typical GGAs and improve the description of correlation effects. SCAN satisfies all 17 exact known constraints that a meta-GGA can satisfy. r2SCAN is a regularized version of SCAN designed to improve numerical stability without hindering accuracy (r2SCAN is meant to obtain similar results to SCAN) [50]. This does so by breaking some of the exact constraints of SCAN.
Although these functionals can simultaneously capture multiple properties such as energy differences between magnetic phases and optimal geometry, SCAN and r2SCAN have been widely reported to overestimate on-site magnetic moments [51, 52, 53, 54]. In the case of 1T-VSe2, we found the DMC on-site magnetic moment of the V atom to be 1.06(2) in Ref. 8. We found the on-site moment to be 1.40 with SCAN and 1.43 with r2SCAN. Nonetheless, due to the success of multi-property predictions with SCAN and r2SCAN with respect to our DMC benchmark calculations, we decided to use SCAN to obtain more complex quantities (that are difficult to obtain with DMC) such as magnetic anisotropy energies. For computationally demanding simulations such as for the Raman active modes where it is not computationally tractable to use SCAN or r2SCAN, we used PBE+ ( = 1 eV) due to its relative qualitative agreement with DMC results (in terms of structural properties).
@@
Figure S1c) depicts the results for the CDW distorted structures with various magnetic orientations. The initial supercells for these distorted structures were obtained from Ref. 26, where the V and Se atoms were initially displaced by 0.12 and 0.18 Å respectively (similarly to Ref. 15) prior to atomic relaxation. For our DFT benchmarking calculations, we adapted a slightly different approach for the distorted CDW structures (in contrast to the undistorted cells). For these supercells, we modified the lattice constant to be a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8) and fixed the cell dimensions. We then allowed the atomic positions (with initially displaced atoms) to relax with each respective DFT functional. Similarly to the undistorted structures in Fig. S1a), there is a massive discrepancy in relative energy between each magnetic orientation (with respect to the NM undistorted cell) of the distorted CDW structures (up to 0.5 eV). We observe that the FM orientation is the lowest energy magnetic state for all functionals depicted in Fig. S1b). In contrast to the undistorted structure in Fig. S1a), the energy differences between the FM and various AFM orientations is much smaller. Our results are in good qualitative agreement with the DFT calculations performed in Ref. 26, but are shifted slightly in energy which can be attributed to fixing the lattice constant to the DMC obtained value. Similar to the undistorted case, we also find excellent qualitative agreement between SCAN/r2SCAN and PBE+ ( = 1 eV) for the energy differences between FM and AFM configurations of distorted structures. To provide more insight on these DFT calculations, we performed DMC simulations for just the FM CDW distorted structure at supercells of 30 and 60 atoms and extrapolated to the thermodynamic limit. For the starting geometry of this structure, we optimized the initially displaced atomic coordinates with r2SCAN and used a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8). This DMC energy difference and uncertainty is shown (with respect to the undistorted NM structure) in Fig. S1c) and Table 1. We find that the DMC energy difference lies between PBE+ ( = 1 eV) and PBE+ ( = 2 eV). Additionally, we find that the SCAN/r2SCAN and LDA+ ( = 3 eV) energy differences are close to the DMC value. More detailed benchmarking data can be found in Table S1, S2, S3, S4 and S5.
@@
From our DMC insight, we went on to estimate the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (). The incorporation of spin-orbit within DMC is a relatively new development, with successful applications in band gap calculations [41] and identifying band inversions in topological materials [55]. For 1T-VSe2, we were mainly focused on the spin-orbit-induced magnetic anisotropy. From our DFT calculations, we know that the magnetic anisotropy energy differences are on the order of fractions of an meV/f.u., while our DMC uncertainties for total energy are on the order of 10 meV/f.u. Since these spin-orbit-induced magnetic anisotropy energy differences are within our DMC uncertainty, we performed collinear DMC calculations (where spins can orient either up or down) for the undistorted FM and AFM orientations (depicted in Fig. S1 and Table 1). For the undistorted structure, we obtain a value of 71(10) meV (no spin-orbit contribution). Taking the insight from our DMC and DFT calculations, where we observed that SCAN quantitatively reproduced our DMC benchmarking results with closer agreement than r2SCAN (see Table 1), we went on to perform spin-orbit SCAN calculations to explicitly obtain anisotropy parameters and (see SI for more detailed information). From these calculations, we obtained a value of -0.57 meV and a value of -0.77 meV. With regards to the distorted-CDW structure, we performed a DMC benchmark calculation for the FM orientation (see Fig. S1 and Table 1). Due to the reasonable agreement with DMC and SCAN and for consistency with the undistorted case, we went on to calculate exchange and anisotropy parameters for the distorted structure using spin-orbit SCAN calculations. In the case of SCAN (since the AFM-A CDW-distorted structure relaxed to the undistorted AFM orientation), the AFM-B2 configuration is the ground state AFM configuration. Using our CDW-FM and CDW-AFM-B2 noncollinear SCAN calculations, we obtained a obtained a value of 21 meV, a value of -0.77 meV, and a value of 0.27 meV (see SI for more details). We went on to use these accurately computed magnetic exchange and anisotropy parameters for our spin Hamiltonian model shown in Eq. 1, where we performed classical Monte Carlo simulations to gain insight on the magnetic phase transitions of the undistorted and distorted structures.